mail A horizontal force of 45 newtons applied to a mass of 9 kilograms is just coefficient to move it if the body is now pulled at an angle of 50 to the horizontal, find the force required to move the body over the horizontal surface
1
Expert's answer
2019-10-17T11:07:51-0400

We have:


"N=mg-F\sin{50}"


"ma=F\cos{50}-\mu (mg-F\sin{50})=0"


The friction coefficient:


"\mu=\frac{F\cos{50}}{mg-F\sin{50}}"

Now:


"ma=F'-\mu mg=0"

"F'=mg\frac{F\cos{50}}{mg-F\sin{50}}"

"F'=(9)(9.8)\frac{45\cos{50}}{(9)(9.8)-45\sin{50}}=47.5\ N"